OpenStudy (anonymous):

A wildlife sanctuary is in the shape of a right triangle as shown. What is the perimeter of the sanctuary? Using complete sentences, explain how you found the perimeter.

OpenStudy (anonymous):

@ganeshie8 can you help please

OpenStudy (anonymous):

@goformit100 can you help please

OpenStudy (anonymous):

can someone tell me if i got this right: Pythagoream helped me get the perimeter of 28

Directrix (directrix):

I don't see how the perimeter could be 28 when two sides of the right triangle are 45 and 53. That is 98 before the third side is added.

OpenStudy (aonz):

First find the length of the 3rd side then add all 3 sides up

OpenStudy (anonymous):

well A(squared)+B(Squared)=C(Squared), so A is 45, B is missing and hyptonuse is C 53

OpenStudy (anonymous):

it's pythagoream theorem

Directrix (directrix):

Did you solve 53^2 = 45^2 x^2 where x is the missing side.

OpenStudy (anonymous):